# BurstGPT: A Real-World Workload Dataset to Optimize LLM Serving Systems

### v3

@@ Abstract
Serving systems for Large Language Models (LLMs) are often optimized to improve quality of service (QoS) and throughput. However, due to the lack of open-sourced LLM serving workloads, these systems are frequently evaluated under unrealistic workload assumptions. Consequently, performance may degrade when these systems are deployed in real-world scenarios.
@@
This work presents BurstGPT, an LLM serving workload with 5.29 million traces from regional Azure OpenAI GPT services over 121 days. BurstGPT captures realistic LLM serving characteristics through detailed tracing of: (1) Concurrency of requests: It traces burstiness variations of requests in Azure OpenAI GPT services, revealing diversified concurrency patterns in different services and model types. (2) Response Lengths of requests: It traces the auto-regressive serving processes of GPT models, showing statistical relations between requests and their responses. (3) Failures of requests: It traces failures of conversation and API services, showing intensive resource needs and limited resource availability of such services in Azure. Details of the characteristics can serve multiple purposes in LLM serving optimizations, such as system evaluation and trace provisioning. In our demo evaluation with BurstGPT, we observe that frequent variations in BurstGPT reveal declines in efficiency, stability, or reliability in realistic LLM serving. We identify that generalization of KV cache management and request scheduling optimization is not guaranteed for different workloads, especially when systems are poorly optimized for unrealistic workloads. We have made the dataset publicly available to encourage further research.
1 Introduction
@@
Large language models (LLMs), particularly Generative Pretrained Transformer (GPT) models, have significantly impacted the AI industry. Notable examples include OpenAI’s ChatGPT [1][2], Google’s Gemini [3], and Meta’s Llama [4]. However, deploying and operating LLM services is costly due to the substantial use of AI accelerators [5]. The high volume of user requests and limited hardware resources further constrain the quality of service (QoS) [5]. Therefore, optimizing serving systems to reduce costs and improve user experience is urgent.
@@
Figure 1: Data Collection and Use Method of BurstGPT. BurstGPT is a real-world workload trace from the Azure OpenAI GPT service. A scaled sample from a period of BurstGPT can be used to optimize serving systems using specific methods, considering realistic concurrency and response patterns. Note that we open-sourced two versions of BurstGPT: a cleaned trace and a raw trace, with failure logs excluded from the cleaned version.
@@
Existing systems primarily optimize system throughput and achieve better QoS. However, these studies often use non-LLM serving workloads in evaluations, derived from synthetic models [6] or non-LLM services like Microsoft Azure Functions (MAF) [7][8]. This work identifies that these unrealistic synthetic workloads do not reflect the workload patterns of auto-regressive LLM serving and limit the effectiveness of evaluating these systems.
@@
In addressing this gap, we introduce a real-world LLM serving workload dataset named BurstGPT. As shown in Figure 1, the traces are collected from the Azure OpenAI GPT Service [9], which serves GPT models on Azure and provides APIs for regional GPT service providers, such as enterprises and campuses, to build customized GPT services. We collected data from one of these regional GPT service providers with over 3,000 users. By deploying a logging engine on this regional service provider, we monitored privacy-independent metrics of each GPT request. In Figure 1, we collected the request start time, request and response token lengths, service type (API or conversational), and LLM type (ChatGPT or GPT-4) over 121 days. The dataset comprises 4.81 million traces of ChatGPT and 0.24 million traces of GPT-4 using API services, and 0.16 million traces of ChatGPT and 0.08 million traces of GPT-4 using conversational services. It also includes failure request while response length is zero.
@@
  • Concurrency of Requests. We identify unique periodic and aperiodic patterns in request concurrency, varying by service and model types. This insight suggests that model-specific and service-specific concurrency evaluations are important for optimizing LLM serving. In our evaluation, (1) concurrency from non-LLM workloads does not accurately reflect system performance as BurstGPT does. (2) Even for LLM workloads, generalization of system optimization across serving types is not guaranteed. We verify that while optimized first-come-first-served scheduling is effective for conversation services, its efficiency can be reduced for API services.
@@
  • Response Lengths of Requests. Longer response lengths lead to higher workloads during LLM serving. The unpredictable nature of response lengths per request introduces uncertainty to system pressure. The statistical study of relationship between request and response lengths in BurstGPT sheds light on future response provisioning, which can help improve system performance.
@@
  • Failures of Requests. We observe a relatively high failure rate in BurstGPT. In our evaluation, we identify that this is primarily due to inefficiencies in KV cache management. Variations in burstiness in BurstGPT lead to memory bottlenecks, causing spikes in failure rates and performance degradation.
2 Preliminary and Motivation
2.1 Limitations of LLM Serving Systems
@@
Each request in LLM serving requires uncertain but large computational due to the model size and auto-regressive nature [10]. To reduce deployment and operational expenses, several specialized frameworks have been developed, including TensorRT-LLM [11], vLLM [10], DeepSpeed [12], and lightLLM [13]. In addition to these frameworks, various memory and compute optimizations have been proposed for efficient decoding [14][15][16][17][18][19][20][21], significantly improving LLM inference performance in deployments. Also, efficient request scheduling is also crucial for optimizing LLM service. For example, Orca [22] introduces iteration-level scheduling, that dynamically adjusts batch size during iterations; FastServe [6] proposes optimized Multilevel-Feedback-Queue (MLFQ) scheduling in the request queue to improve system efficiency.
@@
However, these optimizations have yet to be evaluated on real-world LLM serving workloads. We identify that generalizing serving optimizations from non-LLM or synthetic workloads to LLM workloads can be challenging [23][24] for ensuring serving efficiency, stability, and reliability.
2.2 Towards Workload-aware LLM Serving Systems
@@
Stable workloads lead to consistent system performance across various metrics. However, in real-world, the workload of LLM is diversified due to user, system, and application behaviours. Neglecting any aspect of workload traces during system optimization may yield an incomplete understanding of a framework’s behavior in real-world deployment.
Table 1: Comparison of BurstGPT with Other Traces
Bursty Real-world Synthetic
Workloads BurstGPT MAF1 [7] MAF2 [8] FastServe [6]
Request Lengths
Response Lengths
LLM Concurrency
Failure Requests
Concurrency Patterns
@@
Currently, the assessment of performance in LLM serving frameworks relies on synthetic concurrency [6] or non-LLM concurrency, such as Microsoft Azure Functions (MAF) [7][8]. As concluded in Table 1, these approaches do not reflect real-world LLM workloads due to a lack of empirical data on system and model patterns, as well as user behaviors. For example, the request-per-second (RPS) metric in MAF is much higher (1.64 RPS on average) compared to LLM services (0.019 RPS on average in conversation service and 0.21 RPS on average in API service when using ChatGPT) due to the non-autoregressive and lightweight nature of function services in Azure. However, burstiness in concurrency is much less severe in MAF, likely leading to performance divergence from LLM workloads.
Figure 2: Weekly Periodicity of Conversation Services in BurstGPT.
Figure 3: Weekly Aperiodicity API Services in BurstGPT.
Figure 4: Daily Periodicity Conversation Services in BurstGPT.
Figure 5: Daily Aperiodicity API Services in BurstGPT.
Response and Request Patterns
@@
The computational complexities of prefilling and decoding in LLM systems are $O(s^{2})$ and $O(s)$ relative to the request length $s$, respectively [25]. This relationship between request length and system load significantly differs from that observed in other lightweight cloud functions and synthetic workloads [7][8], highlighting a unique resource occupation of LLM serving. These complexities are also reflected in the request concurrency, failure rates, and latency of each request in LLM serving.
In Section 3, our study comprehensively characterizes patterns of real-world serving workloads. In Section 4, evaluations of serving optimizations on BurstGPT and MAF2[8] highlight the performance divergence between LLM and non-LLM workloads.
3 Patterns in BurstGPT
3.1 Concurrency of Requests
@@
In this study, we measured the number of requests every five or ten minutes over several hours. Each interval was analyzed to determine how frequently the number of requests doubled from the previous five-minute window. For instance, if requests increased from 4 to 8, then from 8 to 16, each doubling was recorded, increasing our ‘two-times count’ respectively. Figure 7 illustrates these doublings in BurstGPT across the monitoring period. It visually represents the frequency of request bursts, showing how often the request load doubled from one time slice to the next. This analysis helps to identify patterns of surges in network demand.
Figure 7 presents the modeled burstiness every twenty minutes in ChatGPT’s conversation and API services during working hours of weekdays. The burstiness is modeled using the Gamma distribution, as supported in [5][6]. We observe that the shape parameter $\alpha$ and the rate parameter $\beta$ of the Gamma distribution vary between conversation and API services. A smaller $\alpha$ indicates a higher coefficient of variation (CV) of the Gamma distribution, meaning the workload is more bursty [6]. In conversation services, the parameter $\alpha$ varies sharply during working hours. In API services, both $\alpha$ and $\beta$ vary sharply during the day. These observations underscore the instability of burstiness in the system.
3.2 Response Lengths of Requests
To identify differences in response length distributions across LLMs, this subsection employs conversational traces in BurstGPT and uses Llama-2-13b-chat[4] for comparison. For Llama-2-13b-chat, we feed 10k real-world prompts to the model for inference and gather the response lengths. The prompts are randomly truncated from the ShareGPT dialogues [26] with alignined distribution to BurstGPT’s conversational traces [27]. This alignment allows for comparing the response length distributions between the two LLMs, facilitating a deeper understanding of their behaviors in responding to conversational prompts.
@@ 3.2 Response Lengths of Requests
In Figure 11, we observe similarities in request distributions and differences in response distributions. Both ChatGPT and Llama-2-13b-chat request distributions adhere to a Zipf distribution [6], characterized by a peak in the frequency of shorter requests. The histogram of ChatGPT’s response tokens reveals a bimodal distribution of length variability. In contrast, Llama-2-13b-chat exhibits a Zipf distribution with a higher frequency of longer requests. These distributions provide insights into the conversational dynamics of request and response lengths for the models. ChatGPT handles longer contexts compared to Llama, reflecting differences in training datasets, model capacities, or intended interactions.
Figure 6: Frequency of Doubling in Request Volume During Workdays in BurstGPT.
Figure 7: Variation of Burstiness in Gamma Distribution in BurstGPT.
@@
3.3 Failures of Requests
We observe a high failure rate in GPT services. Figure 11 shows that ChatGPT has a higher average service failure rate but better stability compared to GPT-4 in both conversational and API interfaces. Overall, GPT-4 has better service reliability over ChatGPT, demonstrating greater consistency and stability with fewer outliers. Additionally, the average failure rate of the conversation service is consistently high, exceeding 5% for ChatGPT, which is significantly higher than that of regular cloud services [28]. We will discuss the reliability problem in detail in Section 4.
Figure 8: Distribution of Request and Response Tokens.
Figure 9: Response Lengths per Request
Figure 10: Response Lengths per Request Bin
Figure 11: Comparison of Statistical Failure Rates in BurstGPT.
4 Demo Use Case: Serving Evaluation
This section presents a sample use case that evaluates LLM serving systems under BurstGPT with streaming, stochastic, and bursty workloads. The evaluation suite is lightweight and modular, requiring minimal code integration for effective deployment. Additionally, it serves as a user-friendly example of how to implement BurstGPT for evaluation and will be integrated and open-sourced within BurstGPT. With our use case, we advocate for using BurstGPT on more serving systems to explore system limitations and optimization opportunities further.
Figure 12: Workload generator overview. It generates simulations of BurstGPT in a burst manner.
4.1 Implementation

### v4

@@ Abstract.
Serving systems for Large Language Models (LLMs) are often optimized to improve quality of service (QoS) and throughput. However, due to the lack of open-source LLM serving workloads, these systems are frequently evaluated under unrealistic workload assumptions. Consequently, performance may degrade when systems are deployed in real-world scenarios.
@@
This work presents BurstGPT, an LLM serving workload with 10.31 million traces from regional Azure OpenAI GPT services over 213 days. BurstGPT captures LLM serving characteristics from user, model and system perspectives: (1) User request concurrency: burstiness variations of requests in Azure OpenAI GPT services, revealing diversified concurrency patterns in different services and model types. (2) User conversation patterns: counts and intervals within conversations for service optimizations. (3) Model response lengths: auto-regressive serving processes of GPT models, showing statistical relations between requests and their responses. (4) System response failures: failures of conversation and API services, showing intensive resource needs and limited availability of LLM services in Azure. The details of the characteristics can serve multiple purposes in LLM serving optimizations, such as system evaluation and trace provisioning. In our demo evaluation with BurstGPT, frequent variations in BurstGPT reveal declines in efficiency, stability, or reliability in realistic LLM serving. We identify that the generalization of KV cache management, scheduling and disaggregation optimizations can be improved under realistic workload evaluations. BurstGPT is publicly available now at https://github.com/HPMLL/BurstGPT and is widely used to develop prototypes of LLM serving frameworks in the industry.
LLM Serving, Workload Trace, Workload Management, Workload Provisioning, System Scheduling
1. Introduction
@@
Large language models (LLMs), particularly Generative Pre-trained Transformer (GPT) models, have significantly impacted the AI industry. Notable examples include OpenAI’s ChatGPT (OpenAI, 2022)(OpenAI, 2023), Google’s Gemini (Team et al., 2024), and Meta’s Llama (Touvron et al., 2023). However, deploying and operating LLM services is costly due to the substantial use of AI accelerators (Li et al., 2023b). The high volume of user requests and limited hardware resources further constrain the quality of service (QoS) (Li et al., 2023b). Therefore, optimizing serving systems to reduce costs and improve user experience is urgent.
@@
Figure 1. Data collection and use method of BurstGPT. BurstGPT is a real-world workload trace from the Azure OpenAI GPT service. A scaled sample from a period of BurstGPT can be used to optimize serving systems using specific methods, considering realistic concurrency and response patterns. Note that we open-sourced two versions of BurstGPT: a cleaned trace and a raw trace, with failure logs excluded from the cleaned version.
@@
Existing systems primarily optimize system throughput and achieve better QoS. However, these studies often use non-LLM serving workloads in evaluations, derived from synthetic models (Wu et al., 2023) or non-LLM services such as Microsoft Azure Functions (MAF) (Shahrad et al., [n. d.])(Zhang et al., 2021). This work identifies that these unrealistic synthetic workloads do not reflect the workload patterns of auto-regressive LLM serving and limit the effectiveness of evaluating these systems.
@@
In addressing this gap, we introduce a real-world LLM serving workload dataset named BurstGPT. As shown in Figure 1, the traces are collected from the Azure OpenAI GPT Service (azu, [n. d.]), which serves GPT models on Azure and provides APIs for regional GPT service providers, such as enterprises and campuses, to build customized GPT services. We collected data from one of these regional GPT service providers with over 3,000 users. By deploying a logging engine on this regional service provider, we monitored privacy-independent metrics of each GPT request. In Figure 1, we collected the request start time, request and response token lengths, service type (API or conversational), and LLM type (ChatGPT, GPT-4 or GPT-4v) for 213 days. The dataset comprises 8.69 million traces of ChatGPT and 0.95 million GPT-4 traces using API services, and 0.30 million ChatGPT traces and 0.16 million traces of GPT-4 using conversational services. It also includes failure requests while the response length is zero.
@@
  • User Request Concurrency. We identify unique patterns in request concurrency and lengths, varying by service types. This insight suggests that service-specific user request pattern evaluations are important for optimizing LLM serving. In our evaluation, (1) concurrency from non-LLM workloads does not accurately reflect system performance as BurstGPT does. (2) Even for LLM workloads, generalization of system optimization across serving types is not guaranteed. For example, we verify that while optimized first-come-first-served scheduling is effective for conversation services, its efficiency can be reduced for API services.
  • User Conversation Patterns. We share analyze traces for the first 2 months in 3.1, 3.3, 3.4 and conversation traces from the last 3 months in 3.2. The detailed analysis of conversation traces can help schedule LLM serving, e.g., KV cache offload strategies.
@@
  • Model Response Lengths. Longer response lengths lead to higher workloads during LLM serving. The unpredictable nature of response lengths per request introduces uncertainty to system pressure. The statistical study of the relationship between request and response lengths in BurstGPT
    sheds light on future response provisioning, which can help improve system performance.
@@
  • System Response Failures. We observe a relatively high failure rate in BurstGPT. In our evaluation, we identify that this is primarily due to inefficiencies in KV cache management. Variations in burstiness in BurstGPT lead to memory bottlenecks, causing spikes in failure rates and performance degradation.
With the dataset, we open-sourced a benchmark suite, BurstGPT-Perf. In our evaluation using BurstGPT-Perf, we first present our insights on using BurstGPT to assess system performance differences between LLM and non-LLM workloads. We also demonstrate use cases of BurstGPT in scheduling policy selection, workload provisioning. Then, we share our insights in prefill-decode(PD) disaggregation in industrial system prototypes. The evaluation results show that BurstGPT can effectively suggest optimizations for LLM serving systems in real-world scenarios.
2. Preliminary and Motivation
Table 1. Comparison of BurstGPT with Other Traces
Serving Real-world Synthetic
Workloads BurstGPT MAF1 (Shahrad et al., [n. d.]) MAF2 (Zhang et al., 2021) FastServe (Wu et al., 2023)
LLM-User Req. Pat.
LLM-User Conv. Pat.
LLM Resp. Len.
Sys. Resp. Fail.
2.1. Limitations of LLM Serving
@@
Each request in LLM serving requires uncertain but large computational due to the model size and auto-regressive nature (Kwon et al., 2023b). To reduce deployment and operational expenses, several specialized frameworks have been developed, including TensorRT-LLM (Neal and Nick, 2023), vLLM (Kwon et al., 2023b), DeepSpeed (Rasley and et.al, 2020), and lightLLM (lig, 2023). In addition to these frameworks, various memory and compute optimizations have been proposed for efficient decoding (Tri et al., 2023)(Hong et al., 2023)(Ge et al., 2023)(Xiao et al., 2023)(Dettmers et al., 2022)(Shazeer, 2019)(Ainslie et al., 2023)(Zhou et al., 2024)(Tang et al., 2024), significantly improving LLM inference performance in deployments. Also, efficient request scheduling is crucial for optimizing LLM service. For example, Orca (Yu et al., 2022) introduces iteration-level scheduling, that dynamically adjusts batch size during iterations; FastServe (Wu et al., 2023) proposes optimized Multilevel-Feedback-Queue (MLFQ) scheduling in the request queue to improve system efficiency.
@@
However, these optimizations have yet to be evaluated on real-world LLM serving workloads. We identify that generalizing serving optimizations from non-LLM or synthetic workloads to LLM workloads can be challenging (Miao et al., 2023)(Miao and et.al, 2023) for ensuring serving efficiency, stability, and reliability.
2.2. Towards Workload-aware LLM Serving
@@
Stable workloads lead to consistent system performance across various metrics. However, in the real world, the workload of LLM is diversified due to user, system, and model behaviors. Neglecting any aspect of workload traces during system optimization may yield an incomplete understanding of a framework’s behavior in real-world deployment.
User Concurrency Patterns
@@
Currently, the assessment of performance in LLM serving frameworks relies on synthetic concurrency (Wu et al., 2023) or non-LLM concurrency, such as Microsoft Azure Functions (MAF) (Shahrad et al., [n. d.])(Zhang et al., 2021). As concluded in Table 1, these approaches do not reflect real-world LLM workloads due to a lack of empirical data on system and model patterns, as well as user behaviors. For example, the request-per-second (RPS) metric in MAF is much higher (1.64 RPS on average) compared to LLM services (0.019 RPS on average in conversation service and 0.21 RPS on average in API service when using ChatGPT) due to the non-autoregressive and lightweight nature of function services in Azure. However, burstiness in concurrency is much less severe in MAF, likely leading to performance divergence from LLM workloads.
Model Patterns
@@
The computational complexities of prefilling and decoding in LLM systems are $O(s^{2})$ and $O(s)$ relative to the request length $s$, respectively (Kwon et al., 2023a). This relationship between request length and system load significantly differs from that observed in other lightweight cloud functions and synthetic workloads (Shahrad et al., [n. d.])(Zhang et al., 2021), highlighting a unique resource occupation of LLM serving. These complexities are also reflected in the request concurrency, failure rates, and latency of each request in LLM serving.
3. Introduction to BurstGPT
Figure 2. Weekly Periodicity of Conversation Services in BurstGPT.
Figure 3. Weekly Aperiodicity API Services in BurstGPT.
Figure 4. Daily Periodicity Conversation Services in BurstGPT.
Figure 5. Daily Aperiodicity API Services in BurstGPT.
3.1. User Request Concurrency
Figure 6. Variation of Burstiness in Gamma Distribution in BurstGPT.
Figure 7. Distribution of Request and Response Tokens.
Figure 8. Response Lengths per Request
@@ 3.1. User Request Concurrency
Figure 6 presents the modeled burstiness every twenty minutes in ChatGPT’s conversation and API services during working hours on weekdays. The burstiness is modeled using the Gamma distribution, as supported in (Li et al., 2023b)(Wu et al., 2023). We observe that the shape parameter $\alpha$ and the rate parameter $\beta$ of the Gamma distribution vary between conversation and API services. A smaller $\alpha$ indicates a higher coefficient of variation (CV) of the Gamma distribution, meaning the workload is more bursty (Wu et al., 2023). In conversation services, the parameter $\alpha$ varies sharply during working hours. In API services, both $\alpha$ and $\beta$ vary sharply during the day. These observations underscore the instability of burstiness in the system.
3.2. User Conversation Patterns
Distribution
Over 35% of conversations end with only one request. The left of Figure 9 shows that the distribution of conversation requests exhibits an exponential decline, with a median of 2 and 75% of conversations with four or fewer requests.
Interval Time
The left of Figure 9 shows the average interval time is not directly related to the number of requests. The P90 interval time increases when the number of requests is less than 5 and then fluctuates after that.
The More Requests, the Longer Intervals
The right of Figure 9 illustrates the number of intervals exceeding a certain duration $c$ in each conversation, which we define as long intervals. We make $c$ equal to 3 different time deltas and the interquartile range (IQR) upper bound. These intervals increase as the number of conversation requests increases.
Figure 9. Count, Average Interval times, and Average Long Intervals as the Number of Requests in a Conversation in BurstGPT.
3.3. Model Response Patterns
To identify differences in response length distributions across LLMs, this subsection employs conversational traces in BurstGPT and uses Llama-2-13b-chat(Touvron et al., 2023) for comparison. For Llama-2-13b-chat, we feed 10k real-world prompts to the model for inference and gather the response lengths. The prompts are randomly truncated from the ShareGPT dialogues (Gudibande et al., 2023) with alignined distribution to BurstGPT’s conversational traces (Zheng et al., 2023). This alignment allows for comparing the response length distributions between the two LLMs, facilitating a deeper understanding of their behaviors in responding to conversational prompts.
@@ 3.3. Model Response Patterns
In Figure 8, we observe similarities in request distributions and differences in response distributions. Both ChatGPT and Llama-2-13b-chat request distributions adhere to a Zipf distribution (Wu et al., 2023), characterized by a peak in the frequency of shorter requests. The histogram of ChatGPT’s response tokens reveals a bimodal distribution of length variability. In contrast, Llama-2-13b-chat exhibits a Zipf distribution with a higher frequency of longer requests. These distributions provide insights into the conversational dynamics of request and response lengths for the models. ChatGPT handles longer contexts compared to Llama, reflecting differences in model capacities or interactions.
@@
Figure 10. Response Lengths per Request Bin
3.4. System Response Failures
We observe a high failure rate in GPT services. Figure 11 shows that ChatGPT has a higher average service failure rate but better stability compared to GPT-4 in both conversational and API interfaces. Overall, GPT-4 has better service reliability over ChatGPT, demonstrating greater consistency and stability with fewer outliers. Additionally, the average failure rate of the conversation service is consistently high, exceeding 5% for ChatGPT, which is significantly higher than that of regular cloud services (Ahsan et al., 2023). We will discuss the reliability problem in detail in Section 4.
Figure 11. Comparison of Statistical Failure Rates in BurstGPT.
4. BurstGPT-Perf: A Benchmark Suite
This section presents our open-sourced benchmark suite: BurstGPT-Perf to evaluate LLM serving systems under BurstGPT with streaming, stochastic, and bursty workloads. The suite is lightweight and modular, requiring minimal code integration for effective deployment. Additionally, it serves as a user-friendly example of how to implement BurstGPT for evaluation and will be integrated and open-sourced within BurstGPT. With our use case, we advocate for using BurstGPT on more serving systems to explore system limitations and optimization opportunities further.
Figure 12. Overview of BurstGPT-Perf. It generates simulations of BurstGPT in a burst manner with two scaling methods: 1. RPS Scaling scales the original BurstGPT data; 2. Modeled Scaling uses Gamma distribution parameters to generate request times and Zipf distribution parameter to generate prompt token lengths.
4.1. Implementation